# Löb-Safe Logics for Reflective Agents

### v1

@@ Abstract
Epistemic and doxastic logics are modal logics for knowledge and belief, and serve as foundational models for rational agents in game theory, philosophy, and computer science. We examine the consequences of modeling agents capable of a certain sort of reflection. Such agents face a formal difficulty due to Löb’s Theorem, called Löb’s Obstacle in the literature. We show how the most popular axiom schemes of epistemic and doxastic logics suffer from Löb’s Obstacle, and present two axiom schemes that that avoid Löb’s Obstacle, which we call Reasonable Löb-Safe Epistemic Doxastic logic ($\mathcal{LSED}^{R}$) and Justified Löb-Safe Epistemic Doxastic logic ($\mathcal{LSED}^{J}$).
@@ 1 Introduction
In Smullyan’s, “Logicians who reason about themselves,” he considers epistemic problems related to undecidability results in mathematics. He identifies “a complete parallelism between logicians who believe propositions and mathematical systems that prove propositions.” In provability logic, the formula $\varphi{\ \Leftrightarrow\ }{\neg}\mathbf{B}_{\mathbf{i}}\,\varphi$ expresses the Gödel proposition, “This proposition is not provable in system $i$.” In a doxastic interpretation, the same formula expresses the reflective belief, “agent $i$ does not believe this proposition.” This means that for any doxastic or epistemic system, if the agents it models are reflective, then it must properly handle the complications that arise from such self-reference.
A reflective agent is one that can form beliefs and knowledge about self-referential sentences and propositions. Such propositions are of the form
$\varphi{\ \Leftrightarrow\ }(\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\psi)$. The reflective proposition “agent $i$ does not believe this proposition” is of that form, where $\psi$ is replaced with $\bot$ in order to formalize it: $\varphi{\ \Leftrightarrow\ }(\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\bot)$, or equivalently, $\varphi{\ \Leftrightarrow\ }({\neg}\mathbf{B}_{\mathbf{i}}\,\varphi)$.
The above uses the $\mathbf{B}_{\mathbf{i}}\,$ operator for belief, but just as easily we could have used $\mathbf{K}_{\mathbf{i}}\,$. The problem for epistemic logic is immediately apparent. Most humans are capable of reasoning about self-referential propositions, so a logic for human knowledge ought to include such propositions. However, with the Truth Axiom of the knowledge operator, this seems to yield inconsistency. Returning to Smullyan’s parallelism, the Truth Axiom, $\mathbf{K}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\varphi$ would translate to provability logic as the axiom “if system $\varphi$ is provable in system $i$, then $\varphi$ is true”. It yields the very same inconsistency that mathematical system $i$ would face if it could prove its own soundness.
The problem is due to Löb in Löb (1955) from the mathematical logic perspective, and Smullyan presents it in Smullyan (1986) from the doxastic logic perspective. It intersects with contemporary research in artificial intelligence foundations, as in Yudkowsky and Hereshoff (2013), which addresses the problem as it pertains to agents being confident in their own conclusions. They have named the problem Löb’s Obstacle, or the Löbian Obstacle. We describe it, and the problem it presents to modal logic of agency.
@@ 2 Löb’s Theorem
The Löb Sentence is sometimes not mentioned as a Condition, because Löb’s Theorem is typically studied in the context of mathematical logic or provability logic, where such self-referential expressiveness is known to exist. We point out, however, that humans are capable of reasoning about self-referential sentences, and any advanced artificial agent will be able to do so, as well. Because systems for representing human-like reasoners should include self-referential sentences and modal fixed points, this condition is satisfied for our concerns. For example, in the foundations of game theory, ideally rational agents can reason about common knowledge among each other, which is itself defined as a modal fixed point. Items (2) and (4) are constants for all normal modal logics. What remains is for an axiom schema to include (3).
@@ 3.4 $\mathcal{KD}$45 Doxastic Logic
3.5 Recent Work by Stalnaker and Bjorndahl
Recent work by Stalnaker and Bjorndahl examine philosophically grounded approaches to combining knowledge and belief modalities.
@@ 3.6 Analysis
Defenders of these logics may wish to deny that these logics can express reflective sentences. However, as we argued earlier, this renders them unsuitable for human-like agents. Furthermore, in many cases, reflective sentences in the form of fixpoints are explicitly included. For example, notions of common knowledge and common belief are frequently added to one of the above logics. These include theorems of the form $\mathbf{C}\varphi\equiv(\varphi{\ \wedge\ }\Lambda_{i\in\mathbb{G}}\mathbf{K}_{\mathbf{i}}\,\mathbf{C}\varphi)$, a straightforwardly reflective sentence, often parsed as “everybody knows that $\varphi$ and everybody knows this sentence”.
@@ 4 Avoiding Löb
An alternative Löb Safe logic uses $\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\langle\mathbf{K}_{i}\rangle\,\mathbf{K}_{\mathbf{i}}\,\varphi$ as a relaxed axiom, which might be called Justified Belief, based on its relating of belief explicitly to evidence of knowledge. We can state this as “$i$ believes $\varphi$ only if it is possible for $i$ given the evidence that $i$ knows $\varphi$. The relationship between “seemingly possible” and “possible given the evidence” is outside of the present scope, but we briefly note that the claims we are here committed to are that “seemingly possible” entails “evidentially possible”, which may intuitively function as a constraint on agents to base perceptions on reasons and evidence, hence our names for Reasonable Belief and Justified Belief. Justified Belief is weaker than Reasonable Belief.
Theorem 4.2 (Reasonable Belief Implies Justified Belief)
@@
This weaker logic, which we call $\mathcal{LSED}^{J}$, for Justified Löb-Safe Epistemic Doxastic logic, is axiomatized as follows.
@@
$\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\langle\mathbf{K}_{i}\rangle\,\mathbf{K}_{\mathbf{i}}\,\varphi$ Justified Belief
@@
Table 5: Logic of $\mathcal{LSED}^{J}$
@@
$\mathcal{LSED}^{J}$ is Löb Safe.
@@
Due to Theorem 4.2, the set of theorems of $\mathcal{LSED}^{J}$ is a subset of those of $\mathcal{LSED}^{R}$. $\mathcal{LSED}^{R}$ is Löb Safe, so $\mathcal{LSED}^{J}$ is as well.
Soundness and Completeness. We once again apply the Sahlqvist-van Benthem Algorithm to generate the frame condition corresponding to Justified Belief.
@@ 6 Conclusion
We responded to this by presenting $\mathcal{LSED}^{R}$ and $\mathcal{LSED}^{J}$, which avoid Löb’s Obstacle while maintaining a defensible model of human-like knowledge and belief, suitable for rational agents interacting with the external world.

### v2

@@ Abstract
Epistemic and doxastic logics are modal logics for knowledge and belief, and serve as foundational models for rational agents in game theory, philosophy, and computer science. We examine the consequences of modeling agents capable of a certain sort of reflection. Such agents face a formal difficulty due to Löb’s Theorem, called Löb’s Obstacle in the literature. We show how the most popular axiom schemes of epistemic and doxastic logics suffer from Löb’s Obstacle, and present two axiom schemes that that avoid Löb’s Obstacle, which we call Reasonable Löb-Safe Epistemic Doxastic logic ($\mathcal{LSED}^{R}$) and Supported Löb-Safe Epistemic Doxastic logic ($\mathcal{LSED}^{S}$).
@@ 1 Introduction
In Smullyan’s, “Logicians who reason about themselves,” he considers epistemic problems related to undecidability results in mathematics. He identifies “a complete parallelism between logicians who believe propositions and mathematical systems that prove propositions.” In provability logic, the formula $\varphi{\ \Leftrightarrow\ }{\neg}\mathbf{B}_{\mathbf{i}}\,\varphi$ expresses the Gödel proposition, “This proposition is not provable in system $i$.” [^2] We ignore subtleties of encodings, here, because they represent the same proposition whether Gödel-encoded or not. In a doxastic interpretation, the same formula expresses the reflective belief, “agent $i$ does not believe this proposition.” This means that for any doxastic or epistemic system, if the agents it models are reflective, then it must properly handle the complications that arise from such self-reference.
A reflective agent is one that can form beliefs and knowledge about self-referential sentences and propositions. A biconditional is used to formalize the self-reference, as in the following: $\varphi{\ \Leftrightarrow\ }(\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\psi)$. The reflective proposition “agent $i$ does not believe this proposition” is of that form, where $\psi$ is replaced with $\bot$ in order to formalize it: $\varphi{\ \Leftrightarrow\ }(\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\bot)$, or equivalently, $\varphi{\ \Leftrightarrow\ }({\neg}\mathbf{B}_{\mathbf{i}}\,\varphi)$.
The above uses the $\mathbf{B}_{\mathbf{i}}\,$ operator for belief, but just as easily we could have used $\mathbf{K}_{\mathbf{i}}\,$. Most humans are capable of reasoning about self-referential propositions, so a logic for human knowledge ought to include such propositions. However, with the Truth Axiom of the knowledge operator, this seems to yield inconsistency. Returning to Smullyan’s parallelism, the Truth Axiom, $\mathbf{K}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\varphi$ would translate to provability logic as the axiom “if formula $\varphi$ is provable in system $i$, then $\varphi$ is true”. This cannot be an axiom in provability logic, as it is for epistemic logic, because it yields the very same inconsistency that mathematical system $i$ would face if it could prove its own soundness, due to Gödel’s Second Incompleteness Theorem.Gödel (1931)
The problem facing epistemic and doxastic logics is due to Löb in Löb (1955) from the mathematical logic perspective, and Smullyan presents it in Smullyan (1986) from the doxastic logic perspective. It intersects with contemporary research in artificial intelligence foundations, as in Yudkowsky and Hereshoff (2013), which addresses the problem as it pertains to agents being confident in their own conclusions. They have named the problem Löb’s Obstacle, or the Löbian Obstacle.
@@ 2 Löb’s Theorem
The Löb Sentence is sometimes not mentioned as a Condition, because Löb’s Theorem is typically studied in the context of mathematical logic or provability logic, where such self-referential expressiveness is known to exist. We point out, however, that humans are capable of reasoning about self-referential sentences, and any advanced artificial agent will be able to do so, as well. Because systems for representing human-like reasoners should include self-referential sentences and modal fixed points, this condition is satisfied for our concerns. For example, in the foundations of game theory, ideally rational agents can reason about common knowledge among each other, which is itself defined as a modal fixed point, as showin in Barwise barwise1988three. Items (2) and (4) are constants for all normal modal logics. What remains is for an axiom schema to include (3).
@@ 3.5 Analysis
Defenders of these logics may wish to deny that these logics can express reflective sentences. However, as we argued earlier, this renders them unsuitable for human-like agents. Furthermore, in many cases, reflective sentences in the form of fixpoints are explicitly included. For example, notions of common knowledge and common belief are frequently added to one of the above logics. These include theorems of the form $\mathbf{C}\varphi\equiv(\varphi{\ \wedge\ }\Lambda_{i\in\mathbb{G}}\mathbf{K}_{\mathbf{i}}\,\mathbf{C}\varphi)$, a straightforwardly reflective sentence, often parsed as “$\varphi$, and everybody knows this sentence”.
@@ 4 Avoiding Löb
An alternative Löb Safe logic uses $\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\langle\mathbf{K}_{i}\rangle\,\mathbf{K}_{\mathbf{i}}\,\varphi$ as a relaxed axiom, which might be called Supported Belief, based on its relating of belief explicitly to evidence of knowledge. We can state this as “$i$ believes $\varphi$ only if some evidence for $i$ supports that $i$ knows $\varphi$. The relationship between “seemingly possible” and “possible given the evidence” is outside of the present scope, but we briefly note that the claims we are here committed to are that “seemingly possible” entails “evidentially possible”, which may intuitively function as a constraint on agents to base perceptions on reasons and evidence, hence our names for Reasonable Belief and Supported Belief. Supported Belief is weaker than Reasonable Belief.
Theorem 4.2 (Reasonable Belief Implies Supported Belief)
@@
This weaker logic, which we call $\mathcal{LSED}^{S}$, for Supported Löb-Safe Epistemic Doxastic logic, is axiomatized as follows.
@@
$\mathbf{B}_{\mathbf{i}}\,\varphi{\ \Rightarrow\ }\langle\mathbf{K}_{i}\rangle\,\mathbf{K}_{\mathbf{i}}\,\varphi$ Supported Belief
@@
Table 5: Logic of $\mathcal{LSED}^{S}$
@@
$\mathcal{LSED}^{S}$ is Löb Safe.
@@
Due to Theorem 4.2, the set of theorems of $\mathcal{LSED}^{S}$ is a subset of those of $\mathcal{LSED}^{R}$. $\mathcal{LSED}^{R}$ is Löb Safe, so $\mathcal{LSED}^{S}$ is as well.
Soundness and Completeness. We once again apply the Sahlqvist-van Benthem Algorithm to generate the frame condition corresponding to Supported Belief.
@@ 6 Conclusion
We responded to this by presenting $\mathcal{LSED}^{R}$ and $\mathcal{LSED}^{S}$, which avoid Löb’s Obstacle while maintaining a defensible model of human-like knowledge and belief, suitable for rational agents interacting with the external world.